# Gx2Mol: De Novo Generation of Hit-like Molecules from Gene Expression Profiles via Deep Learning

### v1

@@ Abstract
$De~{}novo$ generation of hit-like molecules is a challenging task in the drug discovery process. Most methods in previous studies learn the semantics and syntax of molecular structures by analyzing molecular graphs or simplified molecular input line entry system (SMILES) strings; however, they do not take into account the drug responses of the biological systems consisting of genes and proteins. In this study we propose a deep generative model, Gx2Mol, which utilizes gene expression profiles to generate molecular structures with desirable phenotypes for arbitrary target proteins. In the algorithm, a variational autoencoder is employed as a feature extractor to learn the latent feature distribution of the gene expression profiles. Then, a long short-term memory is leveraged as the chemical generator to produce syntactically valid SMILES strings that satisfy the feature conditions of the gene expression profile extracted by the feature extractor. Experimental results and case studies demonstrate that the proposed Gx2Mol model can produce new molecules with potential bioactivities and drug-like properties.
@@ 1 Introduction
In this study, we present a deep generative model, Gx2Mol, to analyze omics data and design new drug structures to overcome the above problems. Specifically, a VAE model is utilized as a feature extractor to learn the low-dimensional features of the gene expression profile data. Then, a long short-term memory (LSTM) [25, 26, 27] model is leveraged as the chemical generator to produce syntactically valid SMILES strings that satisfy the feature conditions of the gene expression profile extracted by the VAE-based feature extractor. The features of the gene expression profiles are involved throughout the training process as conditions for the LSTM model, which can guide the model to generate molecules associated with the target gene expression profile. Our primary contributions are as follows:
@@
The rest of the study is organized as follows. We survey existing work related to various deep learning-based molecular generation methods in section 2. Section 3 presents the proposed Gx2Mol model for $de~{}novo$ molecular generation using gene expression profiles. In section 4, we conduct comprehensive experiments and case studies to compare with the SOTA models to validate the effectiveness of the proposed Gx2Mol model. Finally, we summarize the paper and show the future direction in section 5.
@@ 2.3 Omics Data-driven Molecule Generation
Note that, unlike the previously mentioned $de~{}novo$ molecular generative models, the proposed Gx2Mol aims to generate hit-like molecules that exhibit promising biological activity against specific target proteins or particular diseases, utilizing information from gene expression profiles. Gx2Mol initially extracts biological features from gene expression profiles using a VAE model. Subsequently, these extracted features serve as conditions for the conditional LSTM model, guiding the generation of hit-like molecules.
@@ 3 Proposed Model
Figure 1: Architecture of the proposed Gx2Mol model. (A) A VAE is trained to extract the biological features of gene expression profiles. Here, a VAE encoder attempts to extract the latent feature vector of a gene expression profile, and a VAE decoder attempts to reconstruct the gene expression profile from the latent vector. (B) After the VAE training, the latent vector is utilized as a condition to an LSTM to generate SMILES strings. An extracted latent vector and a vector representation of a start token are concatenated to generate the first atom of a SMILES string. Then, the generated atom and the condition generate the next atom iteratively. This iterative process ends when the defined end token (i.e., $<$EOS$>$) is generated. Finally, all atoms are combined to form a SMILES string. The newly generated SMILES string can be used as a candidate molecule for hit identification to treat diseases.
3.1 Gx2Mol
The proposed Gx2Mol comprises two main components. The feature extractor (i.e., the VAE model) is employed to extract the features of the gene expression profiles, and the generator (i.e., the conditional LSTM model) is used to generate hit-like molecules based on the extracted features of gene expression profiles.
@@ 3.2 Extraction of Biological Features
The architecture of the Gx2Mol model is illustrated in Figure 1. In phase (A), we initiate the process by training a VAE model, extracting essential biological features from gene expression profiles. The encoder network transforms the features of gene expression profiles into a low-dimensional latent space, which is subsequently reconstructed by the decoder. Post the training phase, only the encoder is utilized for subsequent downstream tasks.
Formally, let $\bm{G}=[g_{1},g_{2},\cdots,g_{T}]$ indicate the gene expression profile, where $g_{i}$ represents the $i$-th gene with the maximum gene number of $T$. The VAE model serves as a feature extractor in our Gx2Mol model, tasked with learning a latent feature distribution denoted as $p(z|\bm{G})$. The objective is to align this distribution as closely as possible to the reference distribution $p(z)$, characterized as an isotropic normal distribution. This alignment occurs through the approximation of observed gene expression profiles, while reinforcing the stochastic independence among latent variables. The utilization of the VAE model in this manner facilitates the extraction of meaningful latent features from the input data, as demonstrated in Figure 1 (A). This visualization provides a tangible representation of how the VAE contributes to capturing essential characteristics within the gene expression profiles. High-dimensional gene expression profile reconstruction can be modeled by the integration of the low-dimensional feature space $p(z)$ and conditional distribution $p_{\bm{\theta}}(\bm{G}|z)$ parameterized by ${\bm{\theta}}$:
@@ 3.3 Generation of Hit-like Molecules
During the generation phase, the input to the VAE encoder exclusively comprises gene expression profiles for feature extraction. The resulting extracted features are subsequently employed to steer the process of generating hit-like molecules. Algorithm 1 summarizes the procedure of the proposed Gx2Mol model. Here, sets of the gene expression profiles and SMILES strings are first used to train the feature extractor and chemical generator. In the training phase, the features of gene expression profiles are learned from a VAE-based feature extractor. The extracted features are used as conditions of the LSTM-based molecular generator. In the testing phase, the gene expression profile is employed to generate new hit-like molecules.
Algorithm 1 Procedure for the Gx2Mol model
@@ 4.1 Experimental Setup
Datasets. In this study, we used chemically induced gene expression profiles as training data to train the proposed Gx2Mol model. In addition, we analyzed eight knockdown and two overexpressed target protein perturbed expression profiles to generate hit-like molecules, and disease reversal gene expression profiles as a case study to generate therapeutic molecules.
@@
Dataset splitting and model selection. The dataset was partitioned into distinct sets for training (80%), validation (10%), and testing (10%) to ensure a robust evaluation of our Gx2Mol model. This division allows for effective model training on the training set, tuning of hyperparameters based on the validation set, and unbiased assessment of model performance on the test set. The selection of the optimal model was determined by monitoring the convergence of the total loss function of Gx2Mol during training. Convergence of the loss function indicates stability and optimal performance. This approach ensures the selection of a well-performing model based on its ability to minimize the defined loss and generalize effectively to unseen data.
@@ 4.2 Evaluation Measures
In this study, three statistical indices (validity, uniqueness, and novelty), along with two essential chemical properties (quantitative estimate of drug-likeness (QED) [55] and synthesizability (SA) [56]), and the Tanimoto coefficient [57] were employed to assess hit-like molecules generated by the proposed Gx2Mol model.
@@ 4.3 Gx2Mol Training
Figure 2 shows a comparison of the distribution of fold change values in the gene expression profile of a molecule between the training set and the reconstructed set. Figure 3 shows a comparison of the distribution of fold change values in the average gene expression profile of all molecules between the original set and the reconstructed set. Note that Figure 2 shows the distribution of a gene expression profile of the molecule “C17H25ClN2O3” exposed in the MCF7 cell, whose SMILES representation is denoted as “CCC1=CC(=C(C(=C1O)C(=O)NC[C@@ H]2CCCN2CC)OC)Cl.” The distribution of the original gene expression profiles was similar to that of the reconstructed gene expression profiles acquired using the proposed Gx2Mol model. In other words, the VAE utilized in the proposed Gx2Mol model captures the biological features of the gene expression profiles and successfully reconstructs them into the original distribution.
Note that, in contrast to previous studies like MoFlow [35], FastFlows [36], and DiGress [37], which focused on generating molecules from scratch, our study is specifically oriented towards generating hit-like molecules from gene expression profiles. Furthermore, due to the limited validity of the ConGAN baseline model, which is below 8.5% [23], we restrict our comparison to the proposed Gx2Mol model and the TRIOMPHE baseline model. The statistical results are detailed in Table I, offering a comprehensive comparison with TRIOMPHE. We conducted a comparison between our Gx2Mol model and two variants of the TRIOMPHE model: one based on SELFIES and the other on SMILES-based hit-like molecule generation. Additionally, the LSTM architecture within TRIOMPHE is implemented in two variants, namely unidirectional LSTM (uni-dir) and bidirectional LSTM (bi-dir). In both SELFIES- and SMILES-based molecular generation, the unidirectional LSTM-based TRIOMPHE demonstrated greater uniqueness than its bidirectional LSTM-based model. Furthermore, SMILES-based molecular generation exhibited superior performance compared to SELFIES-based tasks. With the gene expression profiles serving as conditions for Gx2Mol’s LSTM, Gx2Mol generated 1322 molecules. The validity, uniqueness, and novelty of the generated molecules using Gx2Mol are 88.6% (1171 valid molecules), 83.0% (972 unique molecules), and 99.7% (1167 novel molecules). Our Gx2Mol showcases a notable improvement of 10.7% when compared to the best result obtained by the SMILES-based unidirectional TRIOMPHE variant (75.0%). Overall, Gx2Mol model proves to be effective in the task of hit-like molecular generation.
TABLE I: Comparison of key statistical results for the molecules generated by Gx2Mol and baseline models.
@@
Figure 4: Training loss and the ratio of valid molecules generated by the proposed Gx2Mol. The red curve indicates the training loss value of the LSTM with the training epochs. The green curve denotes the ratio of valid molecules generated by the LSTM with the training epochs. Note that the valid molecules are examined by the RDKit tool.
Figure 4 shows the training loss and the ratio of the generated valid molecules of the LSTM in the proposed Gx2Mol model. The loss decreased smoothly over the 300 training epochs and finally converges under 0.1. In contrast, the validity of the molecules generated by the conditional LSTM model gradually increased as training proceeds, with the final validity ratio converging at approximately 90%. Overall, the results indicate that the conditional LSTM model utilized in the proposed Gx2Mol can generate valid molecules effectively.
Figure 5: Violin plots of QED scores for molecules from the training dataset and proposed Gx2Mol.
To further explore the ability of the proposed Gx2Mol model to generate molecules, we also compared the distribution of the QED scores of the molecules generated by Gx2Mol with molecules in the training data. Figure 5 shows that the generated molecules and the original molecules have similar QED distributions. The average QED scores of molecules in the training dataset and molecules generated by Gx2Mol were 0.60 and 0.61, respectively. The violin plots of the QED scores indicate that the proposed Gx2Mol model did not change the potential chemical property characteristics of the training data during the molecular generation process, which demonstrates the LSTM model’s ability to generate molecules effectively.
@@
Figure 7: Top-12 molecular structures with QED scores for molecules generated by Gx2Mol.
Figures 6 and 7 show the top-12 molecular structures with their QED scores for molecules in the training dataset and molecules generated by the proposed Gx2Mol model, respectively. It seems that all of the molecules are chemically valid and exhibit high QED scores.
Furthermore, we evaluated the drug-likeness (QED) scores for the top-k generated molecules using the Gx2Mol model. The results are presented in Table II. The molecules generated by Gx2Mol exhibited QED scores that were higher yet comparable to those of the training data. These findings demonstrate that the Gx2Mol model generated molecules while preserving the inherent QED properties.
@@
Figure 9: Top-12 molecular structures with their SA scores for molecules generated by Gx2Mol.
Similarly, we present the top-12 molecular structures along with their synthesizability (SA) scores for molecules in the training dataset and those generated by Gx2Mol in Figures 8 and 9, respectively. The generated molecular structures indicate that our proposed Gx2Mol can produce valid molecules that are easy to synthesize. Moreover, the SA scores for the top-k generated molecules in Table II demonstrate that Gx2Mol can effectively generate molecules with high SA scores. In other words, the molecules generated by Gx2Mol are confirmed to possess both drug-like characteristics and ease of synthesizability.
4.4 Gx2Mol Generation
Figure 10: Comparison of newly generated molecules from the baseline and Gx2Mol. The 1st column of the table indicates the ten therapeutic target proteins. The 2nd, 4th, and 6th columns indicate known ligands for the corresponding target proteins. The 3rd, 5th, and 7th columns indicate newly generated molecules by ConGAN, TRIOMPHE, and Gx2Mol, which have the highest Tanimoto coefficients with the corresponding known ligands.
@@ 4.4 Gx2Mol Generation
TABLE III: Comparison of structural similarity scores between baseline and Gx2Mol methods.
Therapeutic target protein ConGAN TRIOMPHE Gx2Mol
@@
We conducted experiments on the newly generated molecules by comparing their molecular structures with those of the known ligands. If the newly generated molecules are meaningful, the newly generated molecules should be structurally similar to known ligands of each target protein to some extent. To ensure a fair comparison with the TRIOMPHE baseline, the default sampling number for each gene expression profile of the target protein was set to 1000, consistent with the setting used in TRIOMPHE. Subsequently, we only retained the valid molecules from the 1000 generated samples to calculate structural similarity using Tanimoto coefficients. The results are presented in Table III. ConGAN [23] and TRIOMPHE [24] are the two state-of-the-art (STOA) baseline models that are related to the proposed Gx2Mol model. For the former eight knockdown target proteins, six of the calculated Tanimoto coefficients for the molecules generated by the proposed Gx2Mol model with inhibitory ligands (i.e., AKT1, AKT2, AURKB, CTSK, EGFR, and HDAC1) outperformed the baseline methods. For MTOR and PIK3CCA, the Tanimoto coefficients performed second only to TRIOMPHE. In addition, for both 2SMAD3 and TP53, i.e., the target proteins with gene overexpression perturbations, the Tanimoto coefficients of the generated molecules by the proposed Gx2Mol model were higher than those obtained by the baseline methods.
@@
Figure 10 shows the molecules generated by the baseline and proposed Gx2Mol models. For the 10 target proteins, all generated molecules were structurally similar to the known ligands, compared with the baseline models. In summary, the proposed Gx2Mol model exhibited superior performance in terms of generating hit-like molecules from gene expression profiles via deep learning, and the proposed model outperformed the current SOTA baselines in most metrics.

### v2

# De Novo Generation of Hit-like Molecules from Gene Expression Profiles via Deep Learning

@@ Abstract
$De~{}novo$ generation of hit-like molecules is a challenging task in the drug discovery process. Most methods in previous studies learn the semantics and syntax of molecular structures by analyzing molecular graphs or simplified molecular input line entry system (SMILES) strings; however, they do not take into account the drug responses of the biological systems consisting of genes and proteins. In this study we propose a hybrid neural network, HNNMol, which utilizes gene expression profiles to generate molecular structures with desirable phenotypes for arbitrary target proteins. In the algorithm, a variational autoencoder is employed as a feature extractor to learn the latent feature distribution of the gene expression profiles. Then, a long short-term memory is leveraged as the chemical generator to produce syntactically valid SMILES strings that satisfy the feature conditions of the gene expression profile extracted by the feature extractor. Experimental results and case studies demonstrate that the proposed HVL2Mol model can produce new molecules with potential bioactivities and drug-like properties.
@@ 1 Introduction
In this study, we present a deep generative model, HVL2Mol, to analyze omics data and design new drug structures to overcome the above problems. Specifically, a VAE model is utilized as a feature extractor to learn the low-dimensional features of the gene expression profile data. Then, a long short-term memory (LSTM) [25, 26, 27] model is leveraged as the chemical generator to produce syntactically valid SMILES strings that satisfy the feature conditions of the gene expression profile extracted by the VAE-based feature extractor. The features of the gene expression profiles are involved throughout the training process as conditions for the LSTM model, which can guide the model to generate molecules associated with the target gene expression profile. Our primary contributions are as follows:
@@
The rest of the study is organized as follows. We survey existing work related to various deep learning-based molecular generation methods in section 2. Section 3 presents the proposed HVL2Mol model for $de~{}novo$ molecular generation using gene expression profiles. In section 4, we conduct comprehensive experiments and case studies to compare with the SOTA models to validate the effectiveness of the proposed HVL2Mol model. Finally, we summarize the paper and show the future direction in section 5.
@@ 2.3 Omics Data-driven Molecule Generation
Note that, unlike the previously mentioned $de~{}novo$ molecular generative models, the proposed HVL2Mol aims to generate hit-like molecules that exhibit promising biological activity against specific target proteins or particular diseases, utilizing information from gene expression profiles. HVL2Mol initially extracts biological features from gene expression profiles using a VAE model. Subsequently, these extracted features serve as conditions for the conditional LSTM model, guiding the generation of hit-like molecules.
@@ 3 Proposed Model
Figure 1: Architecture of the proposed HVL2Mol model. (A) A VAE is trained to extract the biological features of gene expression profiles. Here, a VAE encoder attempts to extract the latent feature vector of a gene expression profile, and a VAE decoder attempts to reconstruct the gene expression profile from the latent vector. (B) After the VAE training, the latent vector is utilized as a condition to an LSTM to generate SMILES strings. An extracted latent vector and a vector representation of a start token are concatenated to generate the first atom of a SMILES string. Then, the generated atom and the condition generate the next atom iteratively. This iterative process ends when the defined end token (i.e., $<$EOS$>$) is generated. Finally, all atoms are combined to form a SMILES string. The newly generated SMILES string can be used as a candidate molecule for hit identification to treat diseases.
3.1 HVL2Mol
The proposed HVL2Mol comprises two main components. The feature extractor (i.e., the VAE model) is employed to extract the features of the gene expression profiles, and the generator (i.e., the conditional LSTM model) is used to generate hit-like molecules based on the extracted features of gene expression profiles.
@@ 3.2 Extraction of Biological Features
The architecture of the HVL2Mol model is illustrated in Figure 1. In phase (A), we initiate the process by training a VAE model, extracting essential biological features from gene expression profiles. The encoder network transforms the features of gene expression profiles into a low-dimensional latent space, which is subsequently reconstructed by the decoder. Post the training phase, only the encoder is utilized for subsequent downstream tasks.
Formally, let $\bm{G}=[g_{1},g_{2},\cdots,g_{T}]$ indicate the gene expression profile, where $g_{i}$ represents the $i$-th gene with the maximum gene number of $T$. The VAE model serves as a feature extractor in our HVL2Mol model, tasked with learning a latent feature distribution denoted as $p(z|\bm{G})$. The objective is to align this distribution as closely as possible to the reference distribution $p(z)$, characterized as an isotropic normal distribution. This alignment occurs through the approximation of observed gene expression profiles, while reinforcing the stochastic independence among latent variables. The utilization of the VAE model in this manner facilitates the extraction of meaningful latent features from the input data, as demonstrated in Figure 1 (A). This visualization provides a tangible representation of how the VAE contributes to capturing essential characteristics within the gene expression profiles. High-dimensional gene expression profile reconstruction can be modeled by the integration of the low-dimensional feature space $p(z)$ and conditional distribution $p_{\bm{\theta}}(\bm{G}|z)$ parameterized by ${\bm{\theta}}$:
@@ 3.3 Generation of Hit-like Molecules
During the generation phase, the input to the VAE encoder exclusively comprises gene expression profiles for feature extraction. The resulting extracted features are subsequently employed to steer the process of generating hit-like molecules. Algorithm 1 summarizes the procedure of the proposed HVL2Mol model. Here, sets of the gene expression profiles and SMILES strings are first used to train the feature extractor and chemical generator. In the training phase, the features of gene expression profiles are learned from a VAE-based feature extractor. The extracted features are used as conditions of the LSTM-based molecular generator. In the testing phase, the gene expression profile is employed to generate new hit-like molecules.
Algorithm 1 Procedure for the HVL2Mol model
@@ 4.1 Experimental Setup
Datasets. In this study, we used chemically induced gene expression profiles as training data to train the proposed HVL2Mol model. In addition, we analyzed eight knockdown and two overexpressed target protein perturbed expression profiles to generate hit-like molecules, and disease reversal gene expression profiles as a case study to generate therapeutic molecules.
@@
Dataset splitting and model selection. The dataset was partitioned into distinct sets for training (80%), validation (10%), and testing (10%) to ensure a robust evaluation of our HVL2Mol model. This division allows for effective model training on the training set, tuning of hyperparameters based on the validation set, and unbiased assessment of model performance on the test set. The selection of the optimal model was determined by monitoring the convergence of the total loss function of HVL2Mol during training. Convergence of the loss function indicates stability and optimal performance. This approach ensures the selection of a well-performing model based on its ability to minimize the defined loss and generalize effectively to unseen data.
@@ 4.2 Evaluation Measures
In this study, three statistical indices (validity, uniqueness, and novelty), along with two essential chemical properties (quantitative estimate of drug-likeness (QED) [55] and synthesizability (SA) [56]), and the Tanimoto coefficient [57] were employed to assess hit-like molecules generated by the proposed HVL2Mol model.
@@ 4.3 HVL2Mol Training
Figure 2 shows a comparison of the distribution of fold change values in the gene expression profile of a molecule between the training set and the reconstructed set. Figure 3 shows a comparison of the distribution of fold change values in the average gene expression profile of all molecules between the original set and the reconstructed set. Note that Figure 2 shows the distribution of a gene expression profile of the molecule “C17H25ClN2O3” exposed in the MCF7 cell, whose SMILES representation is denoted as “CCC1=CC(=C(C(=C1O)C(=O)NC[C@@ H]2CCCN2CC)OC)Cl.” The distribution of the original gene expression profiles was similar to that of the reconstructed gene expression profiles acquired using the proposed HVL2Mol model. In other words, the VAE utilized in the proposed HVL2Mol model captures the biological features of the gene expression profiles and successfully reconstructs them into the original distribution.
Note that, in contrast to previous studies like MoFlow [35], FastFlows [36], and DiGress [37], which focused on generating molecules from scratch, our study is specifically oriented towards generating hit-like molecules from gene expression profiles. Furthermore, due to the limited validity of the ConGAN baseline model, which is below 8.5% [23], we restrict our comparison to the proposed HVL2Mol model and the TRIOMPHE baseline model. The statistical results are detailed in Table I, offering a comprehensive comparison with TRIOMPHE. We conducted a comparison between our HVL2Mol model and two variants of the TRIOMPHE model: one based on SELFIES and the other on SMILES-based hit-like molecule generation. Additionally, the LSTM architecture within TRIOMPHE is implemented in two variants, namely unidirectional LSTM (uni-dir) and bidirectional LSTM (bi-dir). In both SELFIES- and SMILES-based molecular generation, the unidirectional LSTM-based TRIOMPHE demonstrated greater uniqueness than its bidirectional LSTM-based model. Furthermore, SMILES-based molecular generation exhibited superior performance compared to SELFIES-based tasks. With the gene expression profiles serving as conditions for HVL2Mol’s LSTM, HVL2Mol generated 1322 molecules. The validity, uniqueness, and novelty of the generated molecules using HVL2Mol are 88.6% (1171 valid molecules), 83.0% (972 unique molecules), and 99.7% (1167 novel molecules). Our HVL2Mol showcases a notable improvement of 10.7% when compared to the best result obtained by the SMILES-based unidirectional TRIOMPHE variant (75.0%). Overall, HVL2Mol model proves to be effective in the task of hit-like molecular generation.
TABLE I: Comparison of key statistical results for the molecules generated by HVL2Mol and baseline models.
@@
Figure 4: Training loss and the ratio of valid molecules generated by the proposed HVL2Mol. The red curve indicates the training loss value of the LSTM with the training epochs. The green curve denotes the ratio of valid molecules generated by the LSTM with the training epochs. Note that the valid molecules are examined by the RDKit tool.
Figure 4 shows the training loss and the ratio of the generated valid molecules of the LSTM in the proposed HVL2Mol model. The loss decreased smoothly over the 300 training epochs and finally converges under 0.1. In contrast, the validity of the molecules generated by the conditional LSTM model gradually increased as training proceeds, with the final validity ratio converging at approximately 90%. Overall, the results indicate that the conditional LSTM model utilized in the proposed HVL2Mol can generate valid molecules effectively.
Figure 5: Violin plots of QED scores for molecules from the training dataset and proposed HVL2Mol.
To further explore the ability of the proposed HVL2Mol model to generate molecules, we also compared the distribution of the QED scores of the molecules generated by HVL2Mol with molecules in the training data. Figure 5 shows that the generated molecules and the original molecules have similar QED distributions. The average QED scores of molecules in the training dataset and molecules generated by HVL2Mol were 0.60 and 0.61, respectively. The violin plots of the QED scores indicate that the proposed HVL2Mol model did not change the potential chemical property characteristics of the training data during the molecular generation process, which demonstrates the LSTM model’s ability to generate molecules effectively.
@@
Figure 7: Top-12 molecular structures with QED scores for molecules generated by HVL2Mol.
Figures 6 and 7 show the top-12 molecular structures with their QED scores for molecules in the training dataset and molecules generated by the proposed HVL2Mol model, respectively. It seems that all of the molecules are chemically valid and exhibit high QED scores.
Furthermore, we evaluated the drug-likeness (QED) scores for the top-k generated molecules using the HVL2Mol model. The results are presented in Table II. The molecules generated by HVL2Mol exhibited QED scores that were higher yet comparable to those of the training data. These findings demonstrate that the HVL2Mol model generated molecules while preserving the inherent QED properties.
@@
Figure 9: Top-12 molecular structures with their SA scores for molecules generated by HVL2Mol.
Similarly, we present the top-12 molecular structures along with their synthesizability (SA) scores for molecules in the training dataset and those generated by HVL2Mol in Figures 8 and 9, respectively. The generated molecular structures indicate that our proposed HVL2Mol can produce valid molecules that are easy to synthesize. Moreover, the SA scores for the top-k generated molecules in Table II demonstrate that HVL2Mol can effectively generate molecules with high SA scores. In other words, the molecules generated by HVL2Mol are confirmed to possess both drug-like characteristics and ease of synthesizability.
4.4 HVL2Mol Generation
Figure 10: Comparison of newly generated molecules from the baseline and HVL2Mol. The 1st column of the table indicates the ten therapeutic target proteins. The 2nd, 4th, and 6th columns indicate known ligands for the corresponding target proteins. The 3rd, 5th, and 7th columns indicate newly generated molecules by ConGAN, TRIOMPHE, and HVL2Mol, which have the highest Tanimoto coefficients with the corresponding known ligands.
@@ 4.4 HVL2Mol Generation
TABLE III: Comparison of structural similarity scores between baseline and HVL2Mol methods.
Therapeutic target protein ConGAN TRIOMPHE HVL2Mol
@@
We conducted experiments on the newly generated molecules by comparing their molecular structures with those of the known ligands. If the newly generated molecules are meaningful, the newly generated molecules should be structurally similar to known ligands of each target protein to some extent. To ensure a fair comparison with the TRIOMPHE baseline, the default sampling number for each gene expression profile of the target protein was set to 1000, consistent with the setting used in TRIOMPHE. Subsequently, we only retained the valid molecules from the 1000 generated samples to calculate structural similarity using Tanimoto coefficients. The results are presented in Table III. ConGAN [23] and TRIOMPHE [24] are the two state-of-the-art (STOA) baseline models that are related to the proposed HVL2Mol model. For the former eight knockdown target proteins, six of the calculated Tanimoto coefficients for the molecules generated by the proposed HVL2Mol model with inhibitory ligands (i.e., AKT1, AKT2, AURKB, CTSK, EGFR, and HDAC1) outperformed the baseline methods. For MTOR and PIK3CCA, the Tanimoto coefficients performed second only to TRIOMPHE. In addition, for both 2SMAD3 and TP53, i.e., the target proteins with gene overexpression perturbations, the Tanimoto coefficients of the generated molecules by the proposed HVL2Mol model were higher than those obtained by the baseline methods.
@@
Figure 10 shows the molecules generated by the baseline and proposed HVL2Mol models. For the 10 target proteins, all generated molecules were structurally similar to the known ligands, compared with the baseline models. In summary, the proposed HVL2Mol model exhibited superior performance in terms of generating hit-like molecules from gene expression profiles via deep learning, and the proposed model outperformed the current SOTA baselines in most metrics.